# Effects of Collaboration on the Performance of Interactive Theme Discovery Systems

### v4

@@ 5 Evaluation
We want to account for the possibility of themes containing a different assortment of documents that otherwise remain relatively semantically similar. To do this, we also measure the semantic similarity between themes using the S-BERT embeddings of the documents assigned to them reimers-gurevych-2019-sentence. We do this in two ways: (1) Centroid similarity, where the similarity of two themes is the similarity between their two centroids, and centroids are calculated by taking the average of all document embeddings within the theme. (2) Group avg. similarity, where the similarity between two themes is the average of pair-wise cosine similarity between all documents in the two themes. As in the Jaccard metric, we report the maximum embedding similarity for each (unique) theme, which we consider to be the “best theme match”, and calculate their averages for comparison across settings.
Table 1 shows the average maximum Jaccard and embedding similarities for each experiment. We find that synchronous groups produce themes with higher average consistency for the topic modeling and relational systems. For the LLM-based system, we did not observe statistical significant differences across collaboration settings. Interestingly, the LLM-based system offered the least amount of intervention/refinement operations for users during the coding process. Whereas the other systems allowed for operations such as splitting and merging topics (topic modeling), and defining relations between concepts (relational approach), the LLM system only allowed users to alter the definition of the theme provided in the prompt. We hypothesize that the more interventions users can make, the more advantage they can take of the deliberation afforded by synchronous coding.
@@
Table 2: Group Avg. Similarity within Annotator Groups. Themes are considered to be more cohesive if intra-theme similarity is high and more distinctive if inter-theme similarity is low. Statistically significant using a two-sample unpaired t-test with $p<0.05$.
@@
Figure 3: Once an annotator has identified themes and they have been propagated the full dataset, we calculate intra-theme similarity by measuring the avg. of the pairwise distances between each document within a theme (left). We calculate inter-theme similarity by measuring the avg. of pairwise distances between each document in a theme and documents assigned to all other themes (right)
@@
Cohesiveness and Distinctiveness. Another dimension for determining the systematicity and clarity of annotation outcomes is by evaluating the similarities and differences between themes within the same code-book. We propose two metrics to measure this: cohesiveness and distinctiveness. A theme is said to be cohesive if its documents are similar to each other (measured by intra-theme similarity) and distinctive if it is dissimilar from documents in other themes within the same code-book (measured by inter-theme similarity). Intuitively, the purpose of grouping documents by theme is to create abstract representations of a dataset, where each theme represents a distinct facet of the data. If themes are not cohesive and distinctive, then it becomes hard to tell which theme a given document should belong to and the code-book falls apart.
@@
Figure 3 shows how to evaluate these metrics for a single annotator (or annotator group). We calculate both the intra-theme similarity and the inter-theme similarity for all the themes in the code-book. Intra-theme similarity is calculated by taking the average of pair-wise similarity between all documents of the same theme. Inter-theme similarity for a given theme is calculated by taking the average pair-wise similarity of documents in that theme with documents in all other themes.
@@
A confounding factor in these measures is that all systems provide broad coverage of documents such that even distantly related documents may be assigned to a theme. To more accurately represent the cohesiveness and distinctiveness of themes in each experiment, we perform the same calculations on a subset comprised of only the top 25% of documents most closely related to each theme. For the relational and LLM-based systems, this top quartile is selected using the distance from the centroid. For the interactive topic model, we use the weights assigned by the model.
@@
Table 2 shows results for both the whole dataset as well as the subset of the documents closest to each theme. Overall, we find that the intra-theme similarities are always higher than inter-theme similarities, which means that themes are at least moderately cohesive and distinctive across the board. For the relational system experiments, we find that the difference between both types of metrics remains stable across collaboration modalities, and that similarities are higher in the synchronous setting. Our findings suggest that themes may be more cohesive but not more distinctive in the synchronous setting.
Results for the interactive topic model and LLM-based systems are nearly identical across collaboration modalities. Given that the operations to characterize themes offered by these two systems are substantially limited (i.e., adding and removing words for the topic model, and refining theme definitions for the LLM system), these results are not surprising. Again, we see that the wider the range of operations afforded to characterize a theme, the more advantage can be taken from synchronous deliberation.
@@
Correctness. Interactive systems allow users to automate large portions of the coding process at the risk of producing inaccurate theme assignments. To estimate how correct the outputs of each system are, we conduct a post-hoc analysis by manually checking a randomly selected sample of $1,200$ document-theme pairs ($200$ per experimental setting). To ensure that our sample is representative of the overall dataset for each experimental setting, we split the data into quartiles based on document relatedness to each theme and select a uniform sample of themes and relatedness scores. As in the previous section, relatedness is calculated using the theme weight distribution for the interactive topic model and distance from the centroid for the the other two systems. To assess reliability, each assignment is evaluated by two annotators, with a third weighing in for tie breaks. We also find that human evaluators have moderate-to-high agreement when assessing system outputs (with an overall Krippendorff’s $\alpha$ of $0.632$), suggesting that we can trust these estimations.
@@
Figure 4 shows the correctness results for each quartile sample per experiment. First, we observe that the relational system is not only the most accurate, but it shows negligible correctness differences in synchronous vs. asynchronous configurations. This an encouraging result, given that this system took the most advantage of synchronous deliberation based on our other metrics of quality. The other two systems showed marked differences, with the topic modeling approach producing more accurate assignments in asynchronous operation, and the LLM producing more accurate assignment in a synchronous paradigm. This observation leads us to believe that interactive systems that introduce less and weaker inductive biases from interaction are more sensitive to choices of configuration – and researchers should be aware of these possibilities when designing experiments. Second, we find that the correctness of different approaches tapers differently, based on the distance of the example from the theme. In other words, as the data samples resemble less like the human labeled ones, some models handle it more robustly than others. In our observation, relational approaches and LLM methods outperform topic models when it comes to assigning themes to distant examples.
@@
Figure 4: Correctness w.r.t. distance from theme.
@@ 6 User Study and Recommendations
We carried out semi-structured interviews with the participants in the study. We focused on trying to understand attitudes towards the task and tools, particularly with respect to the synchronous and asynchronous modalities. Interview questions can be found in App. C. We summarize our main findings below.
Participants found working in teams eases the coding process. Teams of annotators working synchronously described their perceived advantage of working in teams – broader perspectives helped them contextualize diverse tweets easily. The annotators also valued the ability to collaboratively resolve disagreements through discussions. They cited it being easy to "break ties" quickly through conversation. In addition to easing the user experience, we also note that the consistency and cohesiveness of annotations is higher in synchronous settings. Based on this evidence, future work should consider designing systems that explicitly support synchronous deliberation.
Asynchronous coders were more likely to notice inadequacies of the tool they used. Asynchronous annotators directed most of their feedback towards the way the tools functioned. Since the bulk of the annotations were carried in isolation, difficulties navigating the tool were more salient to them than to their synchronous counterparts. Based on these findings, tool designers should focus on improving the asynchronous coding experience.
Annotators noted a lack of control and agency in the topic modeling system. Annotators expressed loss of control and agency when using the topic modeling system fang-2023-user over other systems. One annotator noted that the controls to modify topics were not satisfactory – "the merge process did not offer the ideal amount of control and made it difficult to keep track of the theme groups". Based on these findings, tool designers should consider abstracting users away from the intricacies of the underlying NLP method.
Annotators felt that starting topics often negatively impacted their annotations. Both synchronous and asynchronous annotators were dissatisfied with some of the ways in which tweets belonging to different or opposite themes were placed under the same topic in the topic modeling system. In contrast to the clustering approach used by the other two systems, which appeared to yield cleaner partitions. One of the annotators commented – "Many Anti-Vax and Pro-Vax standpoints use the same words/phrases in their tweets, which the fang-2023-user model groups together despite the stark difference in message between the two.". Some users felt frustrated at being unable to achieve the purity of topic that they hoped for, often stopping when they found that none of their changes were affecting the topics as they intended. However, one annotator did note that the benefit of using a topic modeling system was that there was some amount of theme induction done already - as opposed to the other two systems, where clusters did not have any representative words attached to them. Based on these findings, researchers should investigate ways to effect more change during interactions for topic modeling systems, as well as consider generating some initial explanations to characterize clusters for alternative approaches.
@@
Using LLMs to conduct thematic coding at scale is expensive and less reliable. Despite the recent advances in the reasoning capabilities of LLMs, we observe that their performance at scale is not effective enough to be used for the purpose of propagating annotations. Additionally, running LLMs for the $5$ annotation configurations came with significant time and resource costs. This opens an opportunity for NLP researchers to make LLMs more reliable inductive reasoners, and to come up with prompting strategies that can allow LLMs to reliably classify documents in bulk, especially when working at scale.
Future work developing interactive systems should take into account the amount of control the system offers to the user. Quantitatively, we find that the topic modeling system produced results that were less correct across the board. This matches with our user study where users were unable to disentangle themes that had a high degree of lexical similarity. Despite being able to refine the topic model such that the highest weighted documents matched their defined themes, users had little control over whether or not the rest of the documents in the dataset were faithfully assigned.
@@ 7 Conclusion and Future Work
In this paper, we examined three categories of NLP-assisted qualitative research tools in different collaboration modalities, and conducted inductive and deductive coding on a large dataset of English texts. We designed an evaluation framework that describes the quality of the induced themes and their resulting document assignments under synchronous and asynchronous collaboration. We note that in general, the collaboration modality is a significant factor in determining if the quality of a system’s output. This is particularly true for systems like pacheco-2023-interactive, where users have a wider bandwidth of interaction that can benefit from group consensus, and provide strong inductive biases for deductive coding. We also observe that solutions based on topic modeling, although popular in the data analysis literature, can struggle with inducing cohesive themes and accurate code assignments. Finally, we show that while LLM-based solutions show promise, they pose significant challenges when it comes to coding at scale.
While this study focuses on collaboration modalities, there are numerous other variables that can affect a tool’s efficacy for qualitative coding. We believe that our proposed evaluation framework can be repurposed and expanded to evaluate a wide range of interventions, such as the underlying NLP technology, the interactive interface, the expertise and life experience of the coders, and the type of data being annotated. Through this paper, we hope to inform future work that gives way to more robust measures of an NLP tool’s capabilities as it pertains to qualitative research under realistic settings.
8 Limitations
The study presented in this paper has four main limitations.
(1) While we selected three distinct, representative tools to perform our analysis of synchronous vs. asynchronous settings, the list is of course non-exhaustive. A larger study incorporating more tools could yield additional insights.
(2) As prior work has shown 10.5555/3540261.3540416, automated metrics and human evaluation do not always correlate well for unsupervised or semi-supervised theme induction tasks like topic modeling. Although we did incorporate two types of human evaluation: a post-hoc manual verification (Sec 5) and a user study (Sec 6), both of these evaluations were limited in scope. This is primarily due to how expensive and time consuming it is to perform an exhaustive human evaluation. While we are confident in the findings of our study, a larger study could further strengthen our conclusions.
(3) We used a single dataset in our experiments, which is composed of short English texts. While we believe that this was sufficient to arrive at strong conclusions across systems and collaboration modalities, datasets with different characteristics (e.g., longer documents, non-English language) could pose additional challenges.
@@ 9 Ethical Considerations
Appendix A Topic Model Experimental Settings

### v5

@@ 5.1 Consistency
To account for semantically similar themes with differing document assignments, we also measure semantic consistency using S-BERT document embeddings Reimers and Gurevych (2019). We compute (1) centroid similarity, based on the cosine similarity between theme centroids, and (2) group average similarity, based on the average pairwise similarity across documents in two themes. As with Jaccard, we report maximum similarities per theme and average them for comparison across settings.
Table 1 reports average maximum Jaccard and embedding similarities across experiments. The high variance in Jaccard similarity across datasets, systems, and collaboration settings underscores the importance of semantics-based metrics. For the Covid dataset, synchronous groups produced more consistent themes for the topic modeling and relational systems, while the LLM-based system showed no statistically significant difference across collaboration settings. Results for the Climate dataset were less conclusive, likely due to greater semantic diversity and longer documents, which may encourage synchronous coders to surface more varied themes through discussion; as shown in the next section, these themes are also more cohesive.
Notably, the LLM-based system afforded the fewest opportunities for user intervention during coding. Unlike the other systems, which supported operations such as splitting and merging topics or defining relations between concepts, the LLM system only allowed users to edit theme definitions. We hypothesize that richer intervention mechanisms better enable coders to leverage the deliberation inherent to synchronous collaboration.
@@
Climate All Sync $\mathbf{0.57(0.23)^{*}}$ $0.25(0.08)^{*}$ $\bf 0.44(0.10)^{*}$ $0.30(0.07)^{*}$ $\mathbf{0.39(0.11)^{*}}$ $0.29(0.05)^{*}$
Async $0.50(0.19)^{*}$ $\mathbf{0.24(0.07)}^{*}$ $0.43(0.09)^{*}$ $\bf 0.29(0.07)^{*}$ $0.38(0.08)^{*}$ $\mathbf{0.28(0.05)}^{*}$
Top 25% Sync $\mathbf{0.72(0.21)^{*}}$ $0.26(0.09)^{*}$ $\bf 0.66(0.11)^{*}$ $0.39(0.10)^{*}$ $\mathbf{0.63(0.11)^{*}}$ $0.40(0.09)^{*}$
Async $0.65(0.20)^{*}$ $0.26(0.08)^{*}$ $0.65(0.10)^{*}$ $0.39(0.11)^{*}$ $0.62(0.12)^{*}$ $0.40(0.09)^{*}$
@@
Table 2: Group Avg. Similarity within Coder Groups. Themes are considered to be more cohesive if intra-theme similarity is high and more distinctive if inter-theme similarity is low. Statistically significant using a two-sample unpaired t-test with $p<0.05$.
@@
Figure 3: Once a coder has identified themes and they have been propagated the full dataset, we calculate intra-theme similarity by measuring the avg. of the pairwise distances between each document within a theme (left). We calculate inter-theme similarity by measuring the avg. of pairwise distances between each document in a theme and documents assigned to all other themes (right)
5.2 Cohesiveness and Distinctiveness
@@
Another dimension for determining the systematicity and clarity of coding outcomes is by evaluating the similarities and differences between themes within the same code-book. We propose two metrics to measure this: cohesiveness and distinctiveness. A theme is said to be cohesive if its documents are similar to each other (measured by intra-theme similarity) and distinctive if it is dissimilar from documents in other themes within the same code-book (measured by inter-theme similarity). Intuitively, the purpose of grouping documents by theme is to create abstract representations of a dataset, where each theme represents a distinct facet of the data. If themes are not cohesive and distinctive, then it becomes hard to tell which theme a given document should belong to and the code-book falls apart.
@@
Figure 3 shows how to evaluate these metrics for a single coder (or coder group). We calculate both the intra-theme similarity and the inter-theme similarity for all the themes in the code-book. Intra-theme similarity is calculated by taking the average of pair-wise similarity between all documents of the same theme. Inter-theme similarity for a given theme is calculated by taking the average pair-wise similarity of documents in that theme with documents in all other themes.
@@
A confounding factor in these measures is that all systems provide broad coverage of documents such that even distantly related documents may be assigned to a theme. To more accurately represent the cohesiveness and distinctiveness of the themes in each experiment, we perform the same calculations on a subset comprised of only the top 25% of documents most closely related to each theme. For the relational and LLM-based systems, this top quartile is selected using the distance from the centroid. For the interactive topic model, we use the weights assigned by the model.
@@
Table 2 shows results for both the whole dataset as well as the subset of the documents closest to each theme. Overall, we find that the intra-theme similarities are always higher than inter-theme similarities, which means that themes are at least moderately cohesive and distinctive across the board. For the Covid dataset, we find that themes may be more cohesive but not more distinctive in the synchronous setting especially for the relational system. This may be due to the homogeneity of the dataset, which further explains the uniformity of results in the topic model and LLM experiments.
The results from the Climate dataset present a more compelling finding, where the synchronous experiments uniformly produce much more cohesive themes. When only the top 25% of documents are considered, the increase in cohesiveness is not correlated with a decrease in distinctiveness, unlike in the Covid case. Our results strongly suggest that synchronous collaboration facilitates deeper analysis to uncover clearer themes in the data. This is further supported by the greater distinctiveness in the Climate data experiments: we expect coders to find more distinct themes in a more complex and heterogenous dataset.
5.3 Correctness
@@
Interactive systems allow users to automate large portions of the coding process at the risk of producing inaccurate theme assignments. To estimate how correct the outputs of each system are, we conduct a post-hoc analysis by manually checking a randomly selected sample of $1,200$ document-theme pairs ($200$ per experimental setting). To ensure that our sample is representative of the overall dataset for each experimental setting, we split the data into quartiles based on document relatedness to each theme and select a uniform sample of themes and relatedness scores. As before, relatedness is calculated using the theme weight distribution for the interactive topic model and distance from the centroid for the the other two systems. To assess reliability, each assignment is evaluated by two annotators, with a third weighing in for tie breaks. We also find that human evaluators have moderate-to-high agreement when assessing system outputs (with an overall Krippendorff’s $\alpha$ of $0.632$), suggesting that we can trust these estimations.
@@
Figure 4 shows the correctness results for each quartile sample per experiment in the Covid dataset. First, we observe that the relational system is not only the most accurate, but it shows negligible correctness differences in synchronous vs. asynchronous configurations. This an encouraging result, given that this system took the most advantage of synchronous deliberation based on our other metrics of quality. The other two systems showed marked differences, with the topic modeling approach producing more accurate assignments in asynchronous operation, and the LLM producing more accurate assignment in a synchronous paradigm. This observation leads us to believe that interactive systems that introduce less and weaker inductive biases from interaction are more sensitive to choices of configuration – and researchers should be aware of these possibilities when designing experiments. Second, we find that the correctness of different approaches tapers differently, based on the distance of the example from the theme. In other words, as the data samples resemble less like the human labeled ones, some models handle it more robustly than others. In our observation, relational approaches and LLM methods outperform topic models when it comes to assigning themes to distant examples.
@@
Figure 4: Correctness w.r.t. distance from theme in Covid data experiments.
5.4 Dataset Differences
To what extent do structural and semantic characteristics of a dataset affect the coding process? We carried out a comprehensive analysis across multiple dimensions, including readability, templatic or formulaic patterns, semantic theme ambiguity, and semantic similarity distribution (details in App. A). The higher formulaic content in climate ads (20.1%) may facilitate easy pattern recognition, but could also induce coder fatigue through repetitive content. The greater linguistic complexity (lower Flesch scores, longer sentences) and semantic diversity (lower pairwise similarity) of climate ads suggest higher cognitive load per coding decision.
Conversely, the semantic homogeneity of COVID tweets, evidenced by higher pairwise similarity and centroid proximity, may enable more efficient real-time discussion in synchronous coding settings, as coders share common semantic reference points. The less formulaic nature of the COVID discourse may also sustain coder engagement through content variety. This may explain why the average consistency scores tends to favor synchronous groups for COVID discourse, while for Climate data, asynchronous coders tend to either slightly outperform or are at par with synchronous groups.
These findings suggest that coding protocol selection (synchronous vs. asynchronous) should account for corpus-specific characteristics: semantically diverse, linguistically complex corpora may benefit from asynchronous approaches allowing extended reflection, while homogeneous corpora may be efficiently processed through synchronous collaborative coding.
@@ 6 User Study and Recommendations
We conducted semi-structured interviews to understand the participant experiences with the task and tools, with a focus on synchronous versus asynchronous coding (see App. D for interview script). Several themes emerged.
Synchronous teamwork eased coding and improved outcomes. Participants reported that working synchronously helped them contextualize data, resolve disagreements quickly, and “break ties” through discussion. These experiences align with our quantitative findings showing higher consistency and cohesiveness in synchronous settings, suggesting the value of systems that explicitly support real-time deliberation.
Asynchronous coders were more sensitive to tool limitations. Because they worked largely in isolation, asynchronous participants focused more on usability issues and tool constraints, highlighting the need to improve support for independent coding workflows.
Limited control reduced user trust, particularly in topic modeling. Participants reported a loss of agency when using the topic modeling system, citing insufficient control over operations and difficulty tracking theme evolution – ‘the merge process did not offer the ideal amount of control and made it difficult to keep track of the theme groups”. Initial topics also sometimes conflated opposing themes due to lexical similarity, frustrating users’ attempts to refine results. One participant commented – “Many Anti-Vax and Pro-Vax standpoints use the same words/phrases in their tweets, which the Fang et al. (2023) model groups together despite the stark difference in message between the two.” While some appreciated the model’s initial theme induction, participants desired greater control and clearer explanations of cluster structure. These findings highlight the need for systems to maximize the degree of control afforded to users.
@@
LLM-based coding was costly and unreliable at scale. Despite strong reasoning capabilities, LLMs proved inefficient for large-scale annotation due to high computational costs and inconsistent classification performance. This opens an opportunity for NLP researchers to make LLMs more reliable inductive reasoners and to come up with prompting strategies that can allow LLMs to reliably classify documents in bulk, especially when working at scale.
Overall, our findings suggest that interactive systems should balance automation with user control and provide tailored workflows that support both real-time collaborative deliberation and independent coding at scale.
@@ 7 Conclusion and Future Work
We examined three categories of NLP-assisted qualitative research tools in different collaboration modalities, and conducted inductive and deductive coding on two very different datasets of English texts. We designed an evaluation framework that describes the quality of the induced themes and their resulting document assignments under synchronous and asynchronous collaboration. We find that the collaboration modality is a significant factor in determining if the quality of a system’s output. This is particularly true for systems were users have a wider range of interaction that can benefit from group consensus. We also observe that solutions based on topic modeling, although popular in the data analysis literature, can struggle with inducing cohesive themes and accurate code assignments. Finally, we show that while LLM-based solutions show promise, they pose significant challenges when it comes to coding at scale.
While this study focuses on collaboration modalities, there are numerous other variables that can affect a tool’s efficacy for qualitative coding. We believe that our proposed evaluation framework can be repurposed and expanded to evaluate a wide range of interventions, such as the underlying NLP technology, the interactive interface, the expertise of the coders, and the type of data being annotated. We hope to inform the development of more robust evaluations of NLP tools for qualitative research in realistic settings.
Limitations
The study presented in this paper has two main limitations.
(1) While we selected three distinct, representative tools to perform our analysis of synchronous vs. asynchronous settings, as well as two datasets with distict characteristics, the list is of course non-exhaustive. A larger study incorporating more tools and datasets could yield additional insights.
@@
Appendix A Dataset Analysis
The two corpora exhibit substantial structural differences as seen in Table 3. The climate dataset comprises 5,471 documents with a mean length of 53.1 words, while the COVID dataset contains 85,799 documents averaging 27.5 words. This approximately two-fold difference in document length is statistically significant (Mann-Whitney U, p<0.001, effect size r=-0.353). Climate advertisements also demonstrate greater variability in length (CV=1.47) compared to the more constrained COVID tweets (CV=0.50). Lexical diversity metrics reveal nuanced differences. While type-token ratios are comparable (0.060 vs. 0.051), the Measure of Textual Lexical Diversity (MTLD) indicates higher lexical sophistication in COVID tweets (119.6 vs. 95.6).
Metric Covid Climate
Number of Documents 85799 5471
Words per Document 27.50 53.10
Sentences per Document 2.50 3.40
Type-Token Ratio 0.051 0.060
Measure of Textual
Lexical Diversity
(threshold = 0.72)
119.65 95.58
Table 3: Corpus-level statistics for the Covid and Climate datasets.
Readability analyses indicate that climate advertisements present greater cognitive demands. The Flesch Reading Ease score for climate ads (M=47.4) falls within the “difficult” range, whereas COVID tweets score higher (M=54.9), indicating moderately easier comprehension (p<0.001, r=0.211). Correspondingly, climate ads require higher grade-level reading ability (Flesch-Kincaid Grade: 10.7 vs. 9.2; Gunning Fog Index: 13.2 vs. 11.2). See Table 4.
Metric Covid Climate
Flesch Reading Ease Score 54.85 47.38
Flesch Kincaid Grade 9.18 10.68
Gunning Fog Index 11.20 13.22
Automated Readability Index 12.34 11.85
Table 4: Measures of Linguistic Complexity of the Covid and Climate datasets.
We also performed an analysis over a sample of 5000 documents from both datasets to identify the extent of formulaic or template centric content. DBSCAN clustering on TF-IDF similarity matrices (threshold=0.8) identified 137 template clusters in climate ads, encompassing 20.1% of documents, compared to only 31 clusters (3.7%) in COVID tweets. The proportion of high-similarity document pairs (>=0.8 cosine similarity) is an order of magnitude greater in climate ads (0.059% vs. 0.006%). See Table 5.
Metric Covid Climate
Template Clusters 31 137
Percentage of Documents
in Template Clusters
3.70 20.10
Percentage of High Similarity
Document Pairs
0.006 0.059
Average Cluster Size 5.96 7.34
Largest Cluster Size 31 75
Table 5: Template patterns across Covid and Climate datasets.
Characteristic n-grams reflect domain-specific discourse patterns. Climate ads prominently feature phrases such as “oil natural gas”, “clean energy jobs”, and “fight climate change”, while COVID tweets center on vaccine-related language (“covid 19 vaccine”, “getting covid vaccine”). These patterns suggest that climate advertising employs standardized messaging templates, whereas COVID discourse largely centers around vaccines. See Table 6.
Covid Climate
’covid 19 vaccine’, ’getting covid vaccine’,
’covid vaccine https’, ’got covid vaccine’,
’covid 19 vaccines’
’https bit ly’, ’oil natural gas’,
’clean energy jobs’, ’fight climate change’,
’oil gas industry’
Table 6: Top n-gram phrases in Covid and Climate datasets.
Gaussian Mixture Model analysis with SBERT embeddings (Reimers and Gurevych, 2019) reveals that both corpora exhibit high thematic clarity, with mean maximum cluster probabilities exceeding 0.98. However, COVID tweets demonstrate marginally higher semantic entropy (M=0.042 vs. 0.023), indicating slightly greater theme ambiguity. The proportion of documents spanning multiple clusters (probability difference < 0.2 between top-2 clusters) is higher for COVID tweets (0.66% vs. 0.28%). Silhouette scores are low for both corpora (0.022 vs. 0.018), suggesting that while documents cluster clearly into dominant themes, inter-cluster boundaries are not sharply defined. See Table 7.
Metric Covid Climate
Average Entropy across
Documents
0.042 0.023
Average Confidence in
Primary Cluster Assignment
0.989 0.994
Percentage of Documents
Where Top 2 Clusters are
within 0.2 Probability
0.660 0.280
Silhouette Score 0.022 0.018
Table 7: Semantic theme ambiguity for Covid and Climate datasets.
Mean pairwise cosine similarity in the SBERT embedding space is substantially higher for COVID tweets (0.429 vs. 0.261), indicating greater semantic homogeneity. The inter-quartile range confirms this pattern: COVID tweets exhibit similarity scores between 0.365 – 0.518, while climate ads range from 0.167 – 0.348. Notably, 62.6% of climate ad pairs fall below 0.3 similarity, compared to only 14.1% of COVID tweet pairs. Document-to-centroid similarity further corroborates this finding (0.654 vs. 0.510), demonstrating that COVID tweets cluster more tightly around their corpus centroid. See Table 8.
Metric Covid Climate
Average Pairwise Similarity
between documents
0.43 0.26
25th Percentile Pairwise
Similarity Score
0.37 0.17
75th Percentile Pairwise
Similarity Score
0.52 0.35
Average Similarity of
documents with the centroid
0.65 0.51
Percentage of Pairs with
Similarity > 0.7
0.70 0.20
Percentage of Pairs with
Similarity > 0.3
14.14 62.55
Table 8: Semantic Similarity Distribution for Covid and Climate datasets.
Appendix B Topic Model Experimental Settings